i need to write an equation for the word problem:
To estimate the number of people in Springfield, population 10,000, who have a swimming pool in their backyard, 250 people were interviewed. Of those polled, 89 had a swimming pool. How many people in the city might one expect to have a swimming pool? (Round to the nearest whole number, if necessary.)

Answer by rothauserc(4718)

we solve this using the ratio
250/89 = 10000/x
250x = 10000*89
x = (10000*89) / 250
x = 3560
therefore we expect 3560 swimming pools in the city